sunisround:

Jamie, Pria and Paul can paint a room together in 2 hours. If Pria does the job alone, she can paint the room in 5 hours. If Paul works alone, he can paint the room in 6 hours. If Jamie works alone, how long would it take her to paint the room?

sunisround:

I got (2/x+2)+5/1+6/1=0

sunisround:

but I do not know if i am doing it correct.

Vocaloid:

from what I can remember you calculate the fraction of the job each person does per hour

sunisround:

i am confused.

sunisround:



Vocaloid:

Vocaloid:

I don't remember how to do this one tbh

sunisround:

how in the world do you type this, did you use word?

sunisround:

did you*

Vocaloid:

I just found it on the internet

sunisround:

oh lol

sunisround:

thank you.

Angle:

\[\frac{ together~time }{ Jamie } + \frac{ together~time }{ Pria} + \frac{ together~time }{ Paul } = 1\]

Angle:

it equals 1 because (jamie's part) + (pria's part) + (paul part) = 1 whole

Angle:

in this case (together time) = 2 Pria takes 5 hours Paul takes 6 hours you plus those in and solve for Jamie

Angle:

\[\frac{ 2 }{ 5 } + \frac{ 2 }{ 6 } + \frac{ 2 }{ x } = 1\]

Angle:

if you solve for x from there, you can get your answer :)